OpenStudy (anonymous):

Mark's noraml wage is $880.80 for a 40 hour week. He worked overtime and earned $1233.12 in one week. a) find his normal hourly rate. b) How much extra did he get for overtime? c)how many hours pf overtime did he work if he was paid double time for the overtime worked? i need working out please and i am trying to do them on my own so please just help me

OpenStudy (wasiqss):

hey i told u method, do urself

OpenStudy (anonymous):

wasiqss if you are going to be mean to me dont write or help me then

OpenStudy (anonymous):

a) 880.80 / 40 = 22.01 b) 1233.12 - 880.8 = 352.32 c) 352.32 / (2 * 22.01 ) = 7.97

OpenStudy (wasiqss):

no its a part of rule, nuthin personal,

OpenStudy (anonymous):

wasiquss i just reported u for being mean i am just asking for help

OpenStudy (anonymous):

thankyou for everyone help waskiss ur mean

OpenStudy (wasiqss):

stephaine i bet i wont b kicked or sumthing wud happen to me co im rite